# The magnitude and spectral geometry

### v2

@@ 7.1. Localization of solutions to $A1$ and boundary terms
Since $u_{R}\in H^{-(n+1)/2}(\mathbb{R}^{n})$ is supported in $X$, $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ is supported on $\partial X$. If $n$ is even, it is not clear if $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ is supported on $\partial X$. However, by our result $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ tends to 00 away from $\partial X$ faster than any polynomial as $R\to\infty$. This generalizes [4, Theorem 5] to even dimensions, as relevant for a boundary detection method in data science applications.
@@ 7.2. Taylor expansion at $R=0$
where $Q_{k}(R)u(x)=\frac{1}{k!}\int_{X}\chi(x,y)\mathrm{d}(x,y)^{k}u(y)\mathrm{d}y$ is an elliptic pseudodifferential operator of order $-n-k$ for odd $k$ and smoothing for even $k$ by a similar argument as in Theorem 5.3. We introduce the notation $L_{0}=0$ and
@@ Appendix A Algorithm for computing $c_{j}(X,\mathrm{d})$
    $\quad\quad+\sum_{\begin{subarray}{c}|\alpha^{\prime}|+|\beta^{\prime}|=|\gamma_{l}|,\\ |\alpha^{\prime}|,|\beta^{\prime}|>0\end{subarray}}\frac{\partial_{x^{\prime}}^{\alpha^{\prime}}\varphi(x^{\prime})\partial_{x^{\prime}}^{\beta^{\prime}}\varphi(x^{\prime})}{\alpha^{\prime}!\beta^{\prime}!}D_{\xi^{\prime}}^{\alpha^{\prime}+\beta^{\prime}}\Bigg{]}.$
@@
    Factorize $Q^{\partial}$ near the boundary into terms supported in the upper (+)(+)( + ) and lower ()(-)( - ) half planes
@@ Appendix B Pseudocode in special examples
  • $k$: $n+1-k$ is the degree of term to be determined
@@
  • $N$: dimension of Euclidean space
@@
  • $k$: $n-k$ is the degree of term to be determined

### v3

@@ 7.1. Localization of solutions to $A1$ and boundary terms
Since $u_{R}\in H^{-(n+1)/2}(\mathbb{R}^{n})$ is supported in $X$, $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ is supported on $\partial X$. If $n$ is even, it is not clear if $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ is supported on $\partial X$. However, by our result $u_{R}-\frac{R^{n}}{n!\omega_{n}}$ tends to 0 away from $\partial X$ faster than any polynomial as $R\to\infty$. This generalizes [4, Theorem 5] to even dimensions, as relevant for a boundary detection method in data science applications.
@@ 7.2. Taylor expansion at $R=0$
where $Q_{k}(R)u(x)=\frac{1}{k!}\int_{X}\chi(x,y)\mathrm{d}(x,y)^{k}u(y)\mathrm{d}y$ is an elliptic pseudodifferential operator of order nk-n-k for odd $k$ and smoothing for even $k$ by a similar argument as in Theorem 5.3. We introduce the notation $L_{0}=0$ and
@@ Appendix A Algorithm for computing $c_{j}(X,\mathrm{d})$
    $\qquad\qquad+\sum_{\begin{subarray}{c}|\alpha^{\prime}|+|\beta^{\prime}|=|\gamma_{l}|,\\ |\alpha^{\prime}|,|\beta^{\prime}|>0\end{subarray}}\frac{\partial_{x^{\prime}}^{\alpha^{\prime}}\varphi(x^{\prime})\partial_{x^{\prime}}^{\beta^{\prime}}\varphi(x^{\prime})}{\alpha^{\prime}!\beta^{\prime}!}D_{\xi^{\prime}}^{\alpha^{\prime}+\beta^{\prime}}\Bigg].$
@@
    Factorize $Q^{\partial}$ near the boundary into terms supported in the upper $(+)$ and lower ()(-) half planes
@@ Appendix B Pseudocode in special examples
  • $k$: n+1kn+1-k is the degree of term to be determined
@@
  • NN: dimension of Euclidean space
@@
  • $k$: nkn-k is the degree of term to be determined